OpenStudy (anonymous):

How can x + 2 = -3x + 11 be set up as a system of equations?

OpenStudy (anonymous):

y = x + 13 y = -3x + 11 y = x + 2 y = -3x + 11 y = x + 2 y = -2x + 11 y = 2 y = -3x + 11

OpenStudy (anonymous):

@hartnn